# Graph Out-of-Distribution Generalization via Causal Intervention

### v1

@@ Abstract.
Correspondence author. The work was partly supported by NSFC (92370201, 62222607).
@@ 1. Introduction
Despite great advances in the expressivity and representational power of GNNs, most of existing models focus on improving the performance on in-distribution data, i.e., the testing nodes generated from an identical distribution as the training ones. However, recent evidence (Wu et al., 2022a; Yang et al., 2022; Yu et al., 2023; Yehudai et al., 2021) suggests that GNNs tend to perform unsatisfactorily on out-of-distribution (OOD) data where the data-generating distributions exhibit differences from training observations. We illustrate such an issue through a typical example for node property prediction with GNNs, as shown in Fig. 1. Let us consider a social network where the nodes correspond to users and the goal is to predict whether a user likes playing basketball. In general, if a user’s friends love sports, then the conditional probability for the user liking basketball would be high, which can be treated as a stable (or interchangeably, environment-insensitive) relation from the ego-graph feature (the GNN model actually processes as input of each node) to the label of the target node. Yet, there also exists positive correlation between ”a user’s friends are young” and ”the user likes basketball” on condition that the social network is formed in a university where the marginal probability for ”a user’s friends are young” and ”a user likes playing basketball” are both high. The relation from such an ego-graph feature to the label is unstable (or interchangeably, environment-sensitive), since this correlation does not hold elsewhere like LinkedIn where the marginal distributions for user’s ages and hobbies have considerable diversity. The prediction relying on the latter unstable relation would fail once the environment changes from universities to LinkedIn, which causes distribution shifts. The deficiency of GNNs for OOD generalization urges us to build a shift-robust model for graph representations.
@@
In this paper, we adopt causal analysis from a bottom-up data-generative perspective to investigate the learning behaviors of GNNs under distribution shifts. We reveal that the crux of GNNs’ deficiency for OOD generalization lies in the latent environment confounder that leads to the confounding bias and over-fitting on the environment-sensitive relations. On top of the analysis, we propose a provably effective approach, dubbed as Causal Intervention for Network Data (CaNet), for guiding the GNN to learn stable predictive relations from training data, without prior knowledge of environment labels. We introduce a new learning objective (derived from backdoor adjustment and variational inference) that collaboratively trains an environment estimator and a mixture-of-expert GNN predictor. The former aims to infer pseudo environment labels based on input ego-graphs to partition nodes in the graph into clusters from disparate distributions. The GNN predictor resorts to mixture-of-expert propagation networks dynamically selected by the pseudo environments. The new objective can contribute to alleviating the confounding bias in training data and capturing the environment-insensitive predictive relations that are generalizable across environments.
To evaluate the approach, we conduct experiments on six datasets for node property prediction with various types of distribution shifts. The results manifest that the proposed approach can 1) significantly improve the generalization performance of different GNN models when distribution shifts occur, 2) yield up to $27.4\%$ performance improvements over the state-of-the-art approaches for graph out-of-distribution generalization, and 3) still guarantees competitive performance on in-distribution testing data. We summarize the contributions as follows.
i) We analyze the generalization ability of GNNs under distribution shifts from a causal data-generative perspective, and identify that GNNs trained with maximum likelihood estimation would capture the unstable relations from ego-graph features to labels due to the confounding bias of unobserved environments.
ii) We propose a simple approach for training GNNs under distribution shifts. The model resorts to a novel learning objective that facilitates the GNN to capture environment-insensitive predictive patterns, by means of an environment estimator that infers pseudo environments to eliminate the confounding bias.
iii) We apply our model to various datasets with different types of distribution shifts from training to testing nodes. The results consistently demonstrate the superiority of our model over other graph OOD generalization methods.
@@ 4.1. Causal Analysis for Graph Neural Networks
Figure 2. Structural Causal Models and data pipelines for (a) standard GNNs’ learning process and (b) our proposed approach CaNet’s. The training of common GNNs is affected by the latent confounder $E$ that misguides the model to rely on environment-insensitive correlation between $G$ and $Y$ and leads to unsatisfactory OOD generalization. In contrast, our approach resorts to a new learning objective that essentially cuts off the dependence between $E$ and $G$. This is accomplished with an environment estimator that aims to generate pseudo environments independent of observed data to remove the confounding bias from latent environments.
@@ 4.3. Model Instantiations
Optimization and Algorithm. For model training, we adopt gradient-based optimization for $q_{\phi}$ and $p_{\theta}$ with the objective (5). We assume $p_{0}(D)$ for pseudo environments as a trivial uniform distribution (with equal probabilities for $K$ possible choices) and the loss function induced by (5) can be written as
(11) $-\frac{1}{|\mathcal{V}_{tr}|}\sum_{v\in\mathcal{V}_{tr}}\left[\mathbf{y}_{v}^{\top}\log\hat{\mathbf{y}}_{v}+\frac{\lambda}{L}\sum_{l=1}^{L}\sum_{k=1}^{K}\left[e_{vk}^{(l)}\log\pi_{k}^{(l)}+e_{vk}^{(l)}\log K\right]\right],$
@@
Table 1. Test (mean±$\pm$±standard deviation) Accuracy (%) for citation networks on out-of-distribution (OOD) and in-distribution (ID) data. The distribution shifts are introduced by generating different node features (OOD-Feat) and graph structures (OOD-Struct). OOM indicates out-of-memory error on a GPU with 16GB memory.
@@ 5. Experiments
We apply our model to various node property prediction datasets that involve distribution shifts of different types to evaluate its generalization capability. Overall, we aim to answer the questions:
@@ 5.5. Visualization (R4)
Refer to caption (a) Backbone: GCN Refer to caption (b) Backbone: GAT
@@
Refer to caption (a) Learning curves on Cora Refer to caption (b) Testing accuracy on Arxiv
@@
Refer to caption (a) Arxiv Refer to caption (b) Twitch Refer to caption (c) Arxiv Refer to caption (d) Twitch
@@ A.1. Derivation for Equation 4
We are to establish the equivalence between the objective (5) and $\log p_{\theta}(\hat{Y}|do(G))$. Before the proof, we first introduce two fundamental rules of $do$-calculus (Pearl et al., 2016) which will be used as the building blocks later. Consider a causal directed acyclic graph $\mathcal{A}$ with three nodes: $B$, $D$ and $E$. We denote $\mathcal{A}_{\overline{B}}$ as the intervened causal graph by cutting off all arrows coming into $B$, and $\mathcal{A}_{\underline{B}}$ as the graph by cutting off all arrows going out from $B$. For any interventional distribution compatible with $\mathcal{A}$, the $do$-calculus induces the following two fundamental rules.

### v2

@@ Abstract.
Correspondence author. The work was partly supported by NSFC (92370201, 62222607) and SJTU Trans-med Awards Research (STAR) (20210106).
@@ 1. Introduction
Despite solid advances in the expressivity and representational power of GNNs, most of existing models focus on improving the accuracy on in-distribution data, i.e., the testing nodes generated from an identical distribution as the training ones. However, recent evidence (Wu et al., 2022a; Yang et al., 2022; Yu et al., 2023; Yehudai et al., 2021) suggests that GNNs can perform unsatisfactorily on out-of-distribution (OOD) data where the data-generating distributions exhibit differences from training observations.
@@
We illustrate such an issue through a typical example for node property prediction with GNNs, as shown in Fig. 1. Let us consider a social network where the nodes correspond to users and the goal is to predict whether a user likes playing basketball. In general, if a user’s friends love sports, then the conditional probability for the user liking basketball would be high, which can be treated as a stable (or interchangeably, environment-insensitive) relation from the ego-graph feature (the GNN model actually processes as input of each node) to the label of the target node. Yet, there also exists positive correlation between ”a user’s friends are young” and ”the user likes basketball” on condition that the social network is formed in a university where the marginal probability for ”a user’s friends are young” and ”a user likes playing basketball” are both high. The relation from such an ego-graph feature to the label is unstable (or interchangeably, environment-sensitive), since this correlation does not hold elsewhere like LinkedIn where the marginal distributions for user’s ages and hobbies have considerable diversity. The prediction relying on the latter unstable relation would fail once the environment changes from universities to LinkedIn, which causes distribution shifts. The deficiency of GNNs for OOD generalization urges us to build a shift-robust model for graph representations.
@@
In this paper, we adopt causal analysis from a bottom-up data-generative perspective to investigate the learning behaviors of GNNs under distribution shifts. We reveal that the crux of GNNs’ deficiency for OOD generalization lies in the latent environment confounder that leads to the confounding bias and over-fitting on the environment-sensitive relations. On top of this, we propose a provably effective approach, dubbed as Causal Intervention for Network Data (CaNet), for guiding GNNs to learn stable predictive relations from training data, without prior knowledge of environment labels. We introduce a new learning objective (derived from backdoor adjustment and variational inference) that collaboratively trains an environment estimator and a mixture-of-expert GNN predictor. The former aims to infer pseudo environment labels based on input ego-graphs to partition nodes in the graph into clusters from disparate distributions. The GNN predictor resorts to mixture-of-expert propagation networks dynamically selected by the pseudo environments. The new objective can alleviate the confounding bias in training data and helps to capture the environment-insensitive predictive relations that are generalizable across environments.
To evaluate the approach, we conduct extensive experiments on six graph datasets with various types of distribution shifts. The results manifest that the proposed approach can significantly improve the generalization performance of different GNN models when distribution shifts occur and yield up to $27.4\%$ performance improvements over the state-of-the-arts for graph out-of-distribution generalization. We summarize the contributions below.
$\bullet$   We analyze the generalization ability of GNNs under distribution shifts from a causal data-generative perspective, and identify that GNNs trained with maximum likelihood estimation would capture the unstable relations from ego-graph features to labels due to the confounding bias of unobserved environments.
$\bullet$   We propose a simple yet principled approach for training GNNs under distribution shifts. The model resorts to a novel learning objective that facilitates GNNs to capture environment-insensitive predictive patterns, by means of an environment estimator that infers pseudo environments to eliminate the confounding bias.
$\bullet$   We apply our model to various datasets with different types of distribution shifts from training to testing nodes. The results consistently demonstrate the superiority of our model over other graph OOD generalization methods.
@@ 4.1. Causal Analysis for Graph Neural Networks
Figure 2. Structural Causal Models and data pipelines for (a) standard GNNs’ learning process and (b) our proposed approach CaNet’s. The training of common GNNs is affected by the latent confounder $E$ that misguides the model to rely on environment-insensitive correlation between $G$ and $Y$ and leads to unsatisfactory OOD generalization. In contrast, our approach resorts to a new learning objective that essentially cuts off the dependence between $E$ and $G$.
@@ 4.3. Model Instantiations
Optimization and Algorithm. For model training, we adopt gradient-based optimization for $q_{\phi}$ and $p_{\theta}$ with the objective (5). We assume $p_{0}(E)$ for pseudo environments as a trivial uniform distribution (with equal probabilities for $K$ possible choices) and the loss function induced by (5) can be written as
(11) $\frac{1}{|\mathcal{V}_{tr}|}\sum_{v\in\mathcal{V}_{tr}}\left[-\mathbf{y}_{v}^{\top}\log\hat{\mathbf{y}}_{v}+\frac{\lambda}{L}\sum_{l=1}^{L}\sum_{k=1}^{K}\left[e_{vk}^{(l)}\log\pi_{vk}^{(l)}+e_{vk}^{(l)}\log K\right]\right],$
@@
Table 1. Test (mean±$\pm$±standard deviation) Accuracy (%) for citation networks on out-of-distribution (OOD) and in-distribution (ID) data. OOM indicates out-of-memory error on a GPU with 16GB memory.
@@ 5. Experiments
We apply our model to various datasets to evaluate its generalization capability. Overall, we aim to answer the questions:
@@ 5.5. Visualization (R4)
(a) Backbone: GCN
(b) Backbone: GAT
@@
(a) Learning curves on Cora
(b) Testing accuracy on Arxiv
@@
(a) Arxiv
(b) Twitch
(c) Arxiv
(d) Twitch
@@ A.1. Derivation for Equation 4
We are to derive a tractable objective form for $\log p_{\theta}(\hat{Y}|do(G))$. Before the proof, we first introduce two fundamental rules of $do$-calculus (Pearl et al., 2016) which will be used as the building blocks later. Consider a causal directed acyclic graph $\mathcal{A}$ with three nodes: $B$, $D$ and $E$. We denote $\mathcal{A}_{\overline{B}}$ as the intervened causal graph by cutting off all arrows coming into $B$, and $\mathcal{A}_{\underline{B}}$ as the graph by cutting off all arrows going out from $B$. For any interventional distribution compatible with $\mathcal{A}$, the $do$-calculus induces the following two fundamental rules.